# Single-Image Shadow Removal Using Deep Learning: A Comprehensive Survey

### v1

@@ I Introduction
In recent years, deep learning-based shadow removal methods have demonstrated superior performance, owing to the availability of extensive training data. Figure 1 presents a concise overview of the milestones achieved by deep learning-based shadow removal methods in the past years. In 2017, Qu et al. [24] first proposed a multi-branch fusion framework as the pioneering work to adopt a deep convolutional neural networks (CNNs) [25, 26, 27] to this field. Subsequently, starting from 2018, the advancement of Generative Adversarial Networks (GANs) [28] in the realm of low-level vision has led to widespread adoption of GAN-based networks in shadow removal tasks. These networks are not only effective in addressing amplified artifacts but also enable the development of unsupervised learning methods using unpaired shadow and shadow-free data. During this period, some methods tried to inject constraints, e.g., physical illumination model and mask information, into networks to learn more effective features, by unrolling [29, 30] or recurrent [31] strategies. In the most recent developments, novel techniques such as the transformer mechanism [32, 33, 34, 35] and diffusion model [36, 37, 38, 39, 40] have shown remarkable advancements in shadow removal. The transformer mechanism exploits long-range dependencies within contexts, facilitating more effective extraction of context information from non-shadow regions. Meanwhile, the diffusion model offers a potent diffusive generative prior for generating natural shadow-free images.
@@
Despite the dominance of deep learning in shadow removal research, there is a lack of an in-depth and comprehensive survey on deep learning-based solutions. The existing surveys for shadow removal [41, 42, 43] only cover traditional methods based on earlier publications, but omit the recent works via deep learning approaches. In this paper, we are the first to review and summarize deep learning-based single-image shadow removal methods, aiming to offer a well-structured and comprehensive knowledge base to support and advance the field of shadow removal. The rest of the paper is organized as follows. Section I introduces the shadow formulation model and the problem definition for the shadow removal task. Section II presents a detailed shadow formation model and outlines the various shadow categories, highlighting the different challenges associated with each. Section III provides a detailed survey of deep learning based single-image shadow removal of different learning strategies, including supervised, semi-supervised, unsupervised, and zero-shot learning, as well as summarises the challenges of analysis of shadow removal. Subsequently, Section IV reviews the related technical aspects and explores solutions to the aforementioned challenges through various technical designs and the integration of priors, including network architectures, basic blocks, framework designs, loss functions, datasets, and evaluation metrics. Section V summarizes the quantitative and qualitative comparisons of a number of single-image shadow removal methods among various benchmarks. Section VI presents the related shadow removal applications, including shadow generation and shadow-related attack in practice. Section VII provides detailed discussions of the future research directions on generalized shadow removal and interactive shadow removal. Finally, the paper is concluded in Section VIII.
II Analysis on Shadow Formation
II-A Shadow Formation Model
@@ II-B Shadow Category
III Deep Learning Based Shadow Removal
@@ III-A Problem Definition
III-B General Learning Strategies
@@
Figure 3: Illustration of common deep learning strategies for single-image shadow removal. Details refer to Section III-B.
@@ III-C Challenges and Analysis
IV Technical Review and Discussion
In this section, we review and discuss the technical details and employed priors designed to address the aforementioned challenges in Section III-C from different perspectives, such as network structure, integrating physical illumination models, exploiting shadow detection for removal, and using deep generative models for shadow removal. We then summarize the widely used loss functions, training and testing datasets, and evaluation metrics.
IV-A Network Structure
@@
TABLE I: Summary of key characteristics of notable deep learning-based methods. “Retinex” denotes whether the models are based on the Retinex theory. “Simulated” indicates whether the testing data are simulated in the same manner as the synthetic training data. “Mask” shows whether the models require a mask. “#P” represents the number of trainable parameters.
@@
IV-B Integrating Physical Illumination Model
@@
${\color[rgb]{0,0,0}{\hat{\mathbf{I}}_{sf}}}=\bm{\alpha}\circ\mathbf{I}_{s}+(\mathbf{1}-\bm{\alpha})\circ\mathbf{I}_{lit},$ (6)
@@
IV-C Exploiting Shadow Detection for Removal
@@
    Separate learning: Most shadow removal techniques (refer to the methods that require mask input in Table I) incorporate a pre-trained shadow detection model before the removal stage. This model generates a corresponding shadow mask for the input shadow image, which serves as a conditional input for the shadow removal network. The shadow detection model remains fixed and operates as a pre-processing step for the shadow removal process.
@@
IV-D Deep Generative Model for Shadow Removal
@@
Diffusion-based methods. Recently, numerous diffusion models, including the widely recognized diffusion denoising probability model (DDPM)[37], have garnered significant interest in the field of low-level vision[36, 37, 38, 39]. Guo et al. [30] pioneered the integration of diffusion models into the domain of shadow removal. Subsequently, a series of methods [76, 64, 77] tend to employ the diffusion model as the backbone. While diffusion-based methods have demonstrated their efficacy in producing realistic shadow-free results, they are often accompanied by a significant drawback: the time-consuming nature of the inference process, primarily attributable to the iterative sampling procedures involved. In some applications, such as real-time processing, there may be strict requirements on the latency of inference.
IV-E Loss Function
@@
TABLE II: Summary of shadow datasets. ‘Syn’ represents Synthetic.
Name Number Resolution Format Real/Syn Mask Paired Type
Simulated by Physic Model [17, 93, 94] $\infty$ $\infty$ RGB Syn
Simulated by GAN [51, 56, 48] $\infty$ $\infty$ RGB Syn
UIUC [10] 108 $640\times 425$ RGB Real soft, hard, self
LRSS [17] 137 $5184\times 3456$ raw Real soft
SRD [24] 3,088 $800\times 640$ RGB Real soft, hard
ISTD & ISTD+ [45, 51] 1,870 $640\times 480$ RGB Real hard
USR [56] 2,445 $600\times 450$ RGB Real soft, hard
SBU [95] 5,000 Various RGB Real soft, self, hard
@@
IV-F Training and Testing Datasets
@@
TABLE III: Quantitative comparisons of the shadow removal results on SRD dataset [24] in terms of RMSE, PSNR (in dB), and SSIM. Highlighted in red, yellow, and green cells are the best, second-best, and third-best results, respectively. § indicates that the results are directly quoted from the original paper as their code implementation was not made available.
Learning Method Shadow Region (S) Non-Shadow Region (NS) All Image (ALL)
@@
input 18.96 0.871 36.69 31.47 0.975 4.83 18.19 0.830 14.05 0.1899
TR Guo et al. [10](TPAMI’12) - - 29.89 - - 6.47 - - 12.60 -
SL DeshadowNet [24](CVPR’17) - - 11.78 - - 4.84 - - 6.64 -
DSC [46](TPAMI’19) 30.65 0.960 8.62 31.94 0.965 4.41 27.76 0.903 5.71 0.1141
DHAN [48](AAAI’21) 33.67 0.978 8.94 34.79 0.979 4.80 30.51 0.949 5.67 0.0666
AEF [52](CVPR’21) 32.26 0.966 9.55 31.87 0.945 5.74 28.40 0.893 6.50 0.0894
EMDN [29](AAAI’21) 34.94 0.980 7.44 35.85 0.982 3.74 31.72 0.952 4.79 0.0980
BMNet [54](CVPR’22) 35.05 0.981 6.61 36.02 0.982 3.61 31.69 0.956 4.46 0.0549
ShadowFormer [55](AAAI’23) 35.55 0.982 6.14 36.82 0.983 3.54 32.46 0.957 4.28 0.0572
ShadowDiffusion [30](CVPR’23) 38.72 0.987 4.98 37.78 0.985 3.44 34.73 0.970 3.63 0.0359
Inpaint4Shadow [74](ICCV’23) 36.73 0.985 5.70 36.70 0.985 3.27 33.27 0.967 3.81 0.0475
DeS3 [76](AAAI’24) 37.91 0.986 5.27 37.45 0.984 3.03 34.11 0.968 3.56 0.0338
HomoFormer [78](CVPR’24) 38.81 0.987 4.25 39.45 0.988 2.85 35.37 0.972 3.33 0.0419
UL DC-ShadowNet [58](ICCV’21) 34.00 0.975 7.70 35.53 0.981 3.65 31.53 0.955 4.65 0.1285
SSL ARGAN§ [31](ICCV’19) - - 6.35 - - 4.46 - - 5.31 -
TABLE IV: Quantitative comparisons of the shadow removal results on ISTD+ dataset [45, 51] in terms of RMSE, PSNR (in dB), SSIM, and LPIPS. indicates that the methods use the ground truth shadow mask provided by ISTD+ dataset as input.
Learning Method Shadow Region (S) Non-Shadow Region (NS) All Image (ALL)
@@
input 20.83 0.93 39.01 37.46 0.985 2.40 20.46 0.894 8.40 0.1431
TR Guo et al. [10](TPAMI’12) 26.89 0.960 21.38 35.56 0.976 3.09 25.52 0.925 6.09 0.0966
SL SP+M-Net [51](ICCV’19) 37.60 0.990 5.91 36.02 0.976 2.99 32.94 0.962 3.46 0.0586
Param+M+D-Net [62] 33.09 0.983 9.67 35.35 0.978 2.82 30.15 0.951 3.94 0.0662
AEF [52](CVPR’21) 36.04 0.978 6.55 31.16 0.892 3.77 29.45 0.861 4.23 0.0615
BMNet [54](CVPR’22) 37.87 0.991 5.62 37.51 0.985 2.45 33.98 0.972 2.97 0.0305
SG-ShadowNet [68](ECCV’22) 36.80 0.990 5.93 35.57 0.978 2.92 32.46 0.962 3.41 0.0430
Inpaint4Shadow [74](CVPR’23) 38.10 0.990 6.09 37.66 0.981 2.82 34.16 0.967 3.35 0.0675
ShadowFormer [55](AAAI’23) 39.48 0.992 5.23 38.82 0.983 2.30 35.46 0.971 2.78 0.0260
ShadowDiffusion [30](CVPR’23) 39.69 0.992 4.97 38.89 0.987 2.28 35.67 0.975 2.72 0.0234
HomoFormer [78](CVPR’24) 39.49 0.993 4.73 38.75 0.984 2.23 35.35 0.975 2.64 0.0259
UL DC-ShadowNet [58](ICCV’21) 31.06 0.976 12.62 27.03 0.961 6.82 25.03 0.926 7.77 0.1420
ZSL G2R-ShadowNet [63](CVPR’21) 33.58 0.979 8.82 35.52 0.976 2.89 30.52 0.944 3.86 0.0662
BCDiff [64](ICCV’23) 35.71 0.986 7.61 36.39 0.981 2.66 32.11 0.959 3.47 0.0518
TABLE V: Quantitative comparisons on LRSS dataset [17] and UIUC [10] in terms of RMSE, PSNR (in dB), SSIM, and LPIPS.
Learning Method LRSS UIUC
@@
input 16.43 0.886 15.98 20.85 0.803 13.97
SP+M-Net [51](ICCV’19) 21.77 0.927 11.18 28.23 0.866 7.13
EMDN [29](AAAI’22) 19.26 0.882 15.71 25.56 0.775 13.07
BMNet [54](CVPR’22) 14.19 0.048 50.79 16.86 0.679 45.68
SL Inpaint4Shadow [74](CVPR’23) 19.59 0.805 13.79 27.77 0.849 7.82
ShadowFormer [55](AAAI’23) 27.01 0.957 6.37 28.96 0.874 6.57
ShadowDiffusion [30](CVPR’23) 28.37 0.955 6.01 29.02 0.880 7.05
HomoFormer [78](CVPR’24) 25.86 0.951 6.98 29.08 0.870 6.81
UL DC-ShadowNet [58](ICCV’21) 20.89 0.902 12.55 24.85 0.849 9.51
ZSL G2R-ShadowNet [63](CVPR’21) 20.90 0.901 9.99 27.56 0.858 7.43
BCDiff [64](ICCV’23) 22.13 0.922 10.68 26.81 0.852 7.96
IV-G Evaluation Metrics
@@
V Benchmarking and Empirical Analysis
@@
(c) Mask (DHAN [48])
(d) DSC [46]
(e) DHAN [48]
(f) AEF [52]
(g) EMDN [29]
(h) BMNet [54]
(i) DC-ShadowNet [58]
(j) SG-ShadowNet [68]
(k) Inpaint4Shadow [74]
(l) ShadowDiffusion [30]
(m) DeS3 [76]
(n) HomoFormer [78]
@@
V-A Benchmarking Results
Tables III&IV show the quantitative results on the testsets over ISTD+, and SRD, respectively. When the distribution of testing data is very similar to that of training data, supervised methods achieve better performance. Typically, some recent transformer or diffusion based methods achieved state-of-the-art over ISTD+ and SRD datasets. To further demonstrate the advantages and disadvantages of different methods, Figures 8&9 present the visual examples of the shadow removal results on SRD dataset. For hard shadow removal, the results often exhibit noticeable boundary artifacts due to the difficulty in distinguishing sharp shadow edges from structural edges in the original scene, as illustrated in the lower right corner of Figure 8. It also highlights that in complex scenarios involving multiple shadow types, such as self-shadows and cast shadows, certain shadow regions are often overlooked by shadow detectors. This oversight results in significant residual shadows evident in the outputs of most methods. Inaccurate masks can readily yield incorrect results across all types of shadows. Existing methods have utilized a range of techniques, including joint mask refinement [30], regional classification [76], shadow matte prediction [51], to tackle the issue of mask inaccuracy, which enable better processing of ignored regions.
@@ V-A Benchmarking Results
(c) Mask (DHAN [48])
(d) DSC [46]
(e) DHAN [48]
(f) AEF [52]
(g) EMDN [29]
(h) BMNet [54]
(i) DC-ShadowNet [58]
(j) SG-ShadowNet [68]
(k) Inpaint4Shadow [74]
(l) ShadowDiffusion [30]
(m) DeS3 [76]
(n) HomoFormer [78]
@@
V-B Generalization Capability
To further assess the generalization capabilities of current methods, we utilize the LRSS and UIUC datasets as our evaluation sets. These datasets are specifically chosen because they lack large-scale paired training data, enabling us to rigorously test how well the methods perform in scenarios where there is a distinct distribution difference between training and testing data. For the LRSS dataset, we binarize the residuals of shadow-free images and shadow images to generate mask inputs for methods that require a shadow mask. For the UIUC dataset, we utilize the ground truth masks provided by the dataset. To ensure a fair comparison, we consistently use the ISTD+ training set to train all methods. The performances of most shadow removal methods significantly deteriorate when transferring to testing data that differs markedly from the training data. For instance, models trained on pairs from ISTD+ encounter difficulties in effectively handling shadow images from the soft shadow dataset LRSS, as illustrated in Table V. This is because the ISTD+ dataset predominantly consists of hard shadow types and lacks soft shadow and self-shadow samples. On the contrary, those unsupervised or zero-shot learning methods may have better generalization ability to the unseen scenarios. Some methods, like DC-ShadowNet, considered the classification of shadow types can have a better robustness cross various shadow types. ShadowDiffusion employed the mask refinement as the auxiliary branch to achieve a better robustness to soft shadow processing.
V-C Computational Complexity
In order to provide a comprehensive assessment of deep learning-based shadow removal techniques, we conduct a comparative analysis of the computational complexity of various state-of-the-art methods, as detailed in Table VI, including runtime, training parameters, and FLOPs over 100 images of size $256\times 256$ using one NVIDIA RTX A5000 GPU. Inpaint4Shadow [74] and DC-ShadowNet [58] have the shortest runtimes. BMNet [54] employs a highly efficient architecture based on lightweight invertible blocks, resulting in significantly fewer model parameters and lower FLOPs. Some methods use multiple sub-networks to perform multi-stage processing, which inevitably requires multiples of the parameters and computational load, such as SP+M-Net [51]. Specially, diffusion-based methods, such as ShadowDiffusion [30] and BCDiff [64], require iterative denoising steps. Although multi-stage methods can achieve great performance, they require significantly increased computational complexity and processing time, making them impractical for many applications.
TABLE VI: Quantitative comparisons of the algorithms’ computational efficiency in terms of runtime (in second), number of trainable parameters (#Parameters) (in M), and FLOPs (in G).

### v2

@@ 1 Introduction
In recent years, deep learning-based shadow removal methods [69, 8, 26, 37] have demonstrated superior performance, owing to the availability of extensive training data. Figure 1 presents a concise overview of the milestones achieved by deep learning-based shadow removal methods in the past years. In 2017, Qu et al. [69] first proposed a multi-branch fusion framework as the pioneering work to adopt a deep convolutional neural networks (CNNs) [50, 46, 83] to this field. Subsequently, starting from 2018, the advancement of Generative Adversarial Networks (GANs) [22] in the realm of low-level vision has led to widespread adoption of GAN-based networks in shadow removal tasks. These networks are not only effective in addressing amplified artifacts but also enable the development of unsupervised learning methods using unpaired shadow and shadow-free data. During this period, some methods tried to inject constraints, e.g., physical illumination model and mask information, into networks to learn more effective features, by unrolling [113, 26] or recurrent [8] strategies. In the most recent developments, novel techniques such as the transformer mechanism [82, 9, 57, 53] and diffusion model [79, 32, 42, 71, 60] have shown remarkable advancements in shadow removal. The transformer mechanism exploits long-range dependencies within contexts, facilitating more effective extraction of context information from non-shadow regions. Meanwhile, the diffusion model offers a potent diffusive generative prior for generating natural shadow-free images.
@@
Despite the dominance of deep learning in shadow removal research, there is a lack of an in-depth and comprehensive survey on deep learning-based solutions. The existing surveys for shadow removal [65, 81, 5] only cover traditional methods based on earlier publications, but omit the recent works via deep learning approaches. In this paper, we are the first to review and summarize deep learning-based single-image shadow removal methods, aiming to offer a well-structured and comprehensive knowledge base to support and advance the field of shadow removal. The rest of the paper is organized as follows. Section 1 introduces the shadow formulation model and the problem definition for the shadow removal task. Section 2 presents a detailed shadow formation model and outlines the various shadow categories, highlighting the different challenges associated with each. Section 3 provides a detailed survey of deep learning based single-image shadow removal of different learning strategies, including supervised, semi-supervised, unsupervised, and zero-shot learning, as well as summarises the challenges of analysis of shadow removal. Subsequently, Section 4 reviews the related technical aspects and explores solutions to the aforementioned challenges through various technical designs and the integration of priors, including network architectures, basic blocks, framework designs, loss functions, datasets, and evaluation metrics. Section 5 summarizes the quantitative and qualitative comparisons of a number of single-image shadow removal methods among various benchmarks. Section 6 presents the related shadow removal applications, including shadow generation and shadow-related attack in practice. Section 7 provides detailed discussions of the future research directions on generalized shadow removal and interactive shadow removal. Finally, the paper is concluded in Section 8.
2 Analysis on Shadow Formation
2.1 Shadow Formation Model
@@ 2.2 Shadow Category
3 Deep Learning Based Shadow Removal
@@ 3.1 Problem Definition
3.2 General Learning Strategies
@@
Figure 3: Illustration of common deep learning strategies for single-image shadow removal. Details refer to Section 3.2.
@@ 3.3 Challenges and Analysis
4 Technical Review and Discussion
In this section, we review and discuss the technical details and employed priors designed to address the aforementioned challenges in Section 3.3 from different perspectives, such as network structure, integrating physical illumination models, exploiting shadow detection for removal, and using deep generative models for shadow removal. We then summarize the widely used loss functions, training and testing datasets, and evaluation metrics.
4.1 Network Structure
@@
Table 1: Summary of key characteristics of notable deep learning-based methods. “Physic” denotes whether the methods integrate physical models. “Mask” shows whether the methods require a mask as input for inference.
@@
4.2 Integrating Physical Illumination Model
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\hat{\mathbf{I}}_{sf}}}=\bm{\alpha}\circ\mathbf{I}_{s}+(\mathbf{1}-\bm{\alpha})\circ\mathbf{I}_{lit},$ (6)
@@
4.3 Exploiting Shadow Detection for Removal
@@
    Separate learning: Most shadow removal techniques (refer to the methods that require mask input in Table 1) incorporate a pre-trained shadow detection model before the removal stage. This model generates a corresponding shadow mask for the input shadow image, which serves as a conditional input for the shadow removal network. The shadow detection model remains fixed and operates as a pre-processing step for the shadow removal process.
@@
4.4 Deep Generative Model for Shadow Removal
@@
Diffusion-based methods. Recently, numerous diffusion models, including the widely recognized diffusion denoising probability model (DDPM)[32], have garnered significant interest in the field of low-level vision[79, 32, 42, 71, 93, 91]. Guo et al. [26] pioneered the integration of diffusion models into the domain of shadow removal. Subsequently, a series of methods [39, 27, 62] tend to employ the diffusion model as the backbone. While diffusion-based methods have demonstrated their efficacy in producing realistic shadow-free results, they are often accompanied by a significant drawback: the time-consuming nature of the inference process, primarily attributable to the iterative sampling procedures involved. In some applications, such as real-time processing, there may be strict requirements on the latency of inference.
4.5 Loss Function
@@
Table 2: Summary of shadow datasets. ‘Syn’ represents Synthetic.
Name Number Resolution Format Real/Syn Mask Paired Type Simulated by Physic Model [23, 77, 36] $\infty$ $\infty$ RGB Syn Simulated by GAN [48, 35, 4] $\infty$ $\infty$ RGB Syn UIUC [28] 108 $640\times 425$ RGB Real soft, hard, self LRSS [23] 137 $5184\times 3456$ raw Real soft SRD [69] 3,088 $800\times 640$ RGB Real soft, hard ISTD & ISTD+ [90, 48] 1,870 $640\times 480$ RGB Real hard USR [35] 2,445 $600\times 450$ RGB Real soft, hard SBU [84] 5,000 Various RGB Real soft, self, hard
@@
4.6 Training and Testing Datasets
@@
Table 3: Quantitative comparisons of the shadow removal results on SRD dataset [69] in terms of RMSE, PSNR (in dB), and SSIM. Highlighted in red, yellow, and green cells are the best, second-best, and third-best results, respectively. § indicates that the results are directly quoted from the original paper as their code implementation was not made available.
@@
Table 4: Quantitative comparisons of the shadow removal results on ISTD+ dataset [90, 48] in terms of RMSE, PSNR (in dB), SSIM, and LPIPS. indicates that the methods use the ground truth shadow mask provided by ISTD+ dataset as input.
@@
Table 5: Quantitative comparisons on LRSS dataset [23] and UIUC [28] in terms of RMSE, PSNR (in dB), SSIM, and LPIPS.
@@
4.7 Evaluation Metrics
@@
5 Benchmarking and Empirical Analysis
@@
(c) Mask (DHAN)
(d) DSC
(e) DHAN
(f) AEF
(g) EMDN
(h) BMNet
(i) DC-ShadowNet
(j) SG-ShadowNet
(k) Inpaint4Shadow
(l) ShadowDiffusion
(m) DeS3
(n) HomoFormer
@@
5.1 Benchmarking Results
Tables 3&4 show the quantitative results on the testsets over ISTD+, and SRD, respectively. When the distribution of testing data is very similar to that of training data, supervised methods achieve better performance. Typically, some recent transformer or diffusion based methods achieved state-of-the-art over ISTD+ and SRD datasets. To further demonstrate the advantages and disadvantages of different methods, Figures 8&9 present the visual examples of the shadow removal results on SRD dataset. For hard shadow removal, the results often exhibit noticeable boundary artifacts due to the difficulty in distinguishing sharp shadow edges from structural edges in the original scene, as illustrated in the lower right corner of Figure 8. It also highlights that in complex scenarios involving multiple shadow types, such as self-shadows and cast shadows, certain shadow regions are often overlooked by shadow detectors. This oversight results in significant residual shadows evident in the outputs of most methods. Inaccurate masks can readily yield incorrect results across all types of shadows. Existing methods have utilized a range of techniques, including joint mask refinement [26], regional classification [39], shadow matte prediction [48], to tackle the issue of mask inaccuracy, which enable better processing of ignored regions.
@@ 5.1 Benchmarking Results
(c) Mask (DHAN)
(d) DSC
(e) DHAN
(f) AEF
(g) EMDN
(h) BMNet
(i) DC-ShadowNet
(j) SG-ShadowNet
(k) Inpaint4Shadow
(l) ShadowDiffusion
(m) DeS3
(n) HomoFormer
@@
5.2 Generalization Capability
To further assess the generalization capabilities of current methods, we utilize the LRSS and UIUC datasets as our evaluation sets. These datasets are specifically chosen because they lack large-scale paired training data, enabling us to rigorously test how well the methods perform in scenarios where there is a distinct distribution difference between training and testing data. For the LRSS dataset, we binarize the residuals of shadow-free images and shadow images to generate mask inputs for methods that require a shadow mask. For the UIUC dataset, we utilize the ground truth masks provided by the dataset. To ensure a fair comparison, we consistently use the ISTD+ training set to train all methods. The performances of most shadow removal methods significantly deteriorate when transferring to testing data that differs markedly from the training data. For instance, models trained on pairs from ISTD+ encounter difficulties in effectively handling shadow images from the soft shadow dataset LRSS, as illustrated in Table 5. This is because the ISTD+ dataset predominantly consists of hard shadow types and lacks soft shadow and self-shadow samples. On the contrary, those unsupervised or zero-shot learning methods may have better generalization ability to the unseen scenarios. Some methods, like DC-ShadowNet, considered the classification of shadow types can have a better robustness cross various shadow types. ShadowDiffusion employed the mask refinement as the auxiliary branch to achieve a better robustness to soft shadow processing.
5.3 Computational Complexity
In order to provide a comprehensive assessment of deep learning-based shadow removal techniques, we conduct a comparative analysis of the computational complexity of various state-of-the-art methods, as detailed in Table 6, including runtime, training parameters, and FLOPs over 100 images of size $256\times 256$ using one NVIDIA RTX A5000 GPU. Inpaint4Shadow [52] and DC-ShadowNet [38] have the shortest runtimes. BMNet [112] employs a highly efficient architecture based on lightweight invertible blocks, resulting in significantly fewer model parameters and lower FLOPs. Some methods use multiple sub-networks to perform multi-stage processing, which inevitably requires multiples of the parameters and computational load, such as SP+M-Net [48]. Specially, diffusion-based methods, such as ShadowDiffusion [26] and BCDiff [27], require iterative denoising steps. Although multi-stage methods can achieve great performance, they require significantly increased computational complexity and processing time, making them impractical for many applications.
Table 6: Quantitative comparisons of the algorithms’ computational efficiency in terms of runtime (in second), number of trainable parameters (#Parameters) (in M), and FLOPs (in G).